OpenStudy (anonymous):

Determine an equation of the line with the given slope m that passes through point m=0, (-1,3)

OpenStudy (anonymous):

Are you familiar with point slope formula? It makes doing this very easy

myininaya (myininaya):

or you can use my favorite slope intercept form

OpenStudy (anonymous):

I am very unsure of what forms to use with this one. I do know the formula when you have 2 sets of points.

OpenStudy (anonymous):

But for point slope you need two points...

OpenStudy (anonymous):

When you have a point, and a slope. Use point slope. \[y-y_0 = m(x-x_0)\] Where m in the slope, and \(x_0,y_0\) are the x and y of the point.

OpenStudy (anonymous):

Try it for the one you have here.

OpenStudy (anonymous):

y-3=0(x-(-1))

OpenStudy (anonymous):

Yep. Now simplify.

OpenStudy (anonymous):

You still worken on it?

OpenStudy (anonymous):

It should be very easy since 0 times anything is 0.

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

y-3 = 0

OpenStudy (anonymous):

So y = ?

OpenStudy (anonymous):

Okay, that is what I have on my paper. The reason I just put 0, is because all the examples show y-3=0...........then they show y=0

OpenStudy (anonymous):

No, y = 3

OpenStudy (anonymous):

That is why it's so confusing. The book is really bad.

OpenStudy (anonymous):

So the line you described would be horizontal and maintain a constant value of 3 for any value of x.

OpenStudy (anonymous):

So I literally write the equation y-3=0 since it wants an equation; not just an answer. Does that sound right?

OpenStudy (anonymous):

You do need to write an equation. y = 3 is an equation.

OpenStudy (anonymous):

Oh wow.... I really do appreciate your help. You make it so easy but this book; oh my goodness.

OpenStudy (anonymous):

You want your equation in the form of y = something

OpenStudy (anonymous):

Where something has an x or a constant term, or both.

OpenStudy (anonymous):

okay Thank s alot